# A Panchromatic View of Late-time Shock Power in the Type II Supernova 2023ixf

### v2

@@ 4 Conclusions
Following the transition to a regime dominated by shock powered emission, the observed UV/optical/NIR spectra are dominated by emission from the CDS and inner ejecta. For example, H shows broad, “boxy” emission from a CDS with velocity of – this structure also being seen in UV (e.g., Mg ii , Ly-) and NIR (e.g., He i, Pa-) emission lines. Significant blue-red asymmetry observed in H, combined with luminous near- and mid-IR emission, indicates dust formation in the inner ejecta beginning at days and then in the CDS by at least days.
We quantify the relative contributions from absorbed radioactive decay power and shock powered emission in SN 2023ixf by constructing m SEDs at days. In the first epoch, we find erg -1 and erg -1, which results in an emergent shock powered emission of erg -1. Then, at days, we find erg -1, erg -1, and erg -1.
@@ Appendix A Additional X-ray Analysis
2025-01-26T17:10:18 | 60393.6 | 619 | HST | STIS/FUV MAMA | 0.11-0.17 | (Bostroem et al., 2024) |
2025-01-26T14:27:10 | 60393.5 | 619 | HST | STIS/NUV MAMA | 0.16-0.32 | (Bostroem et al., 2024) |

### v3

@@ 4 Conclusions
At late-time phases, H shows broad, “boxy” emission from a CDS with velocity of – this structure also being seen in UV (e.g., Mg ii , Ly-) and NIR (e.g., He i, Pa-) emission lines. We interpret these observations as a transition to a regime dominated by shock powered emission where the UV/optical/NIR spectra are dominated by emission from the CDS and inner ejecta. Significant blue-red asymmetry observed in H, combined with luminous near- and mid-IR emission, indicates dust formation in the inner ejecta beginning at days and then in the CDS by at least days.
We quantify the relative contributions from absorbed radioactive decay power and shock powered emission in SN 2023ixf by constructing m SEDs at days. In the first epoch, we find erg s-1 and erg s-1, which results in an emergent shock powered emission of erg s-1. Then, at days, we find erg s-1, erg s-1, and erg s-1.
@@ 5 Acknowledgments
Some/all of the data presented in this paper were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi.org/10.17909/aj20-5b11 (catalog https://doi.org/10.17909/aj20-5b11). STScI is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS5–26555. Support to MAST for these data is provided by the NASA Office of Space Science via grant NAG5–7584 and by other grants and contracts.
@@ Appendix A Additional X-ray Analysis
| 2025-01-26T17:10:18 | 60393.6 | 619 | HST | STIS/FUV MAMA | 0.11-0.17 | (Bostroem et al., 2024b) |
| 2025-01-26T14:27:10 | 60393.5 | 619 | HST | STIS/NUV MAMA | 0.16-0.32 | (Bostroem et al., 2024b) |